OpenStudy (anonymous):

Select an ordered pair from the choices below that is a solution to the following system of inequalities: 2y < -5 + x -x + y < 7 (5, 0) (-1, -4) (3, -1) (2, 5)

OpenStudy (anonymous):

can you help me with this one @gypsy1274

OpenStudy (anonymous):

Well, you can do this two ways... You can plug in each value for x and y (I think that's cheating and a failing of multiple choice questions) or you can solve this system of inequalities using substitution or addition/subtraction.

OpenStudy (anonymous):

First, let me see if I understand the question: \(2y<-5+x\) and \(-x+y<7\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Here is my demonstration of systems of equations by substitution: inequalities are treated the same way, except that the inequality will reverse when multiplying or dividing by a negative number. \(2x – 3y = –2\) and \(4x + y = 24\) I'm going to choose the second equation and solve for y. \(4x+y=24\) First, isolate y by subtracting both sides by \(4x\): \(y = 24 − 4x\) If y had a coefficient(other than 1) you would divide both sides by that number. It doesn't so we are done with this step. Now, Substitute \(24−4x\) for y in the other equation: \(2x − 3(24 − 4x) = −2\) And solve this equation using Order of Operations. Parenthesis: \(2x − 72 + 12x = −2\) Combine like terms by adding 72 to each side then combine the x terms: \(14x = 70\) Divide both sides by 14 to isolate the x: \(x = 5\) Now, substitute 5 into each equation and solve for y: (hint: you should get the same answer on each equation - this is how you know you are correct.) \(2x – 3y = –2\) and \(4x + y = 24\) \(2(5) – 3y = –2\) and \(4(5) + y = 24\) \(10 – 3y = –2\) and \(20 + y = 24\) \(−3y = −12\) and \( y = 4\) \(y = 4\)

OpenStudy (anonymous):

? thats alot to take in

OpenStudy (anonymous):

Yes it is. One step at a time (and that is a good point, I should number the steps.) Pick one equation and one variable to solve for.

OpenStudy (anonymous):

I would suggest a variable that has no coefficient. It just makes the calculations easier, but it really doesn't matter.

OpenStudy (anonymous):

my mom helped me on this one sorry this was a question i already did i forgot iorgot i did it oops silly me thank you tho sorry

OpenStudy (anonymous):

Mistakes make us human. And now I know to break up my explanation into smaller bits. Thanks for pointing that out. :-)

OpenStudy (anonymous):

your welcome :)